George invested some money at 5% interest. George also invested $219more than 44 times that amount at 10%. How much is invested at each rate if George receives $922.80 in interest after one year? (Round to two decimal places if necessary.)

------- amount invested at 5%
-------- amount invested at 10%

Answer by greenestamps(13200)



x = amount at 5%
44x+219 = amount at 10%
.05(x) = interest at 5%
.10(44x+219) = interest at 10%

The total interest is $922.80:

.05(x) + .10(44x+219) = 922.80

Solve....